# Riemannian Patch Assignment Gradient Flows

### v1

@@ Abstract
Keywords:
assignment flows Riemannian gradient flows statistical manifolds information geometry.
@@ 1 Introduction
Assignment flows [3] denote a class of dynamical systems evolving on statistical manifolds which serve as models for ‘neural ODEs’ on graphs obtained by geometric flow integration. A basic instance [16] are systems of the form $\dot{W}=R_{W}[\Omega W]$ (see Eq. 12 for the general form) which are Riemannian gradient flows with respect to a potential, parametrized by a spatial interaction matrix $\Omega(t)$, comprising a non-local graph Laplacian and a term steering the flow towards discrete decisions (label assignment). The time-variant parameters $\Omega(t)$ can be learned from data.
By vectorization (see Section 3.1 for details), this basic flow equation takes the form $\dot{w}=R^{\mathfrak{v}}_{w}[(\Omega\otimes I_{c})w]$ which suggests to generalize the parametrization to $\dot{w}=R^{\mathfrak{v}}_{w}[(\Omega\otimes\Omega_{c})w]$, in order to take additionally into account the interaction between $[c]=\{1,\dotsc,c\}$ labels, to be jointly assigned to data observed on a graph via the assignment flow $w(t)$. In fact, this formulation is general enough to cover a broad range of multi-population and multi-game dynamics [5].
@@
Related work. There is a vast literature on image denoising and restoration using continuous patches, either in a classical way by enforcing structure sparsity [9, 13, 19] or more recently by, e.g., Gaussian mixture priors [1, 11]. By contrast, our approach exploits labeled patches, where each label represents an equivalence class of continuous signals, for the structured prediction of labelings of metric data observed on a graph.
@@ 2 Preliminaries
We introduce few definitions from information geometry required in connection with assignment flows [2], [3], [4, Chapter 2]. Given a set $\mathcal{C}$ of category labels, we denote by $\mathcal{S}_{c}=\{w\in\mathbb{R}_{>0}:\langle\mathbb{1}_{c},w\rangle=1\}$ the relative interior of the probability simplex over $c=|\mathcal{C}|$ labels. This space forms a smooth manifold with trivial tangent bundle $T\mathcal{S}_{c}=\mathcal{S}_{c}\times T_{0}$ and tangent space at any $w\in\mathcal{S}_{c}$ given by $T_{0}:=T_{w}\mathcal{S}_{c}=\{v\in\mathbb{R}^{c}:\langle\mathbb{1}_{c},v\rangle=0\}$. $(\mathcal{S}_{c},g^{\mathcal{S}_{c}})$ becomes a Riemannian manifold when equipped with the Fisher-Rao metric
@@
Assignment flows [3] constitute an approach for determining a labeling function (2), based on the evolution of label assignment vectors as integral curves of vector fields on the assignment manifold, governed by a coupled system of replicator equations
@@
The function $F:\mathcal{S}_{c}^{n}\to\mathbb{R}^{n\times c}$ parametrizing the flow is often referred to as fitness (or payoff) function in evolutionary game theory and population dynamics, where analyses of individual (uncoupled) replicator equations have provided foundational insights into evolutionary stability and strategic dynamical games [10, 14]. For a more detailed exposition of assignment flows, we refer to [17].
@@
Geometric integration of such flows can be performed numerically in a stable way using any of the methods worked out by [21]. Under suitable conditions on the fitness function [22], these schemes guarantee convergence $\lim_{t\to\infty}W_{i}(t)=:e_{\ell_{\mathcal{V}}(i)},\;i\in\mathcal{V}$ to some labeling function (2), and they are amenable to learning parameters of the affinity function $F(W)$ from data [12, 20, 6, 8].
@@ 3.1 Flow Definition
Lemma 1 (maximizing patch consistency)
@@
Proof
(16) results from (15) by taking into account the adjacency relation of $\mathcal{G}_{\mathcal{V}}$ given by $A^{h}_{\mathcal{V}}+A^{v}_{\mathcal{V}}=A_{\mathcal{V}}$, due to (1) and (7). The inner product with the first term on the right hand side of (15) reads
@@
Using the row-stacking vectorization map $\operatorname{vec}_{r}(ABC)=(A\otimes C^{\top})\operatorname{vec}_{r}(B)$ [18], we define the vectorized objective function
@@ 3.2 Independence of Graph Orientation
Proof (sketch)
Let $\widetilde{\mathcal{G}}_{\mathcal{V}}=(\mathcal{V},\widetilde{\mathcal{E}}_{\mathcal{V}})$ denote the transpose of $\mathcal{G}_{\mathcal{V}}$, that is the graph with the same vertex set $\mathcal{V}$ but with reversed edge orientations: $ij\in\mathcal{E}_{\mathcal{V}}\;\Leftrightarrow\;ji\in\widetilde{\mathcal{E}}_{\mathcal{V}},\;\text{for all }i,j\in\mathcal{V}$. Since the orientation affects the patch template adjacency matrix by (8), the resulting patch adjacency graph $\widetilde{\mathcal{G}}_{\mathcal{D}}$ becomes transposed as well. As a result, following Section 3.1, the corresponding objective then reads $\widetilde{J}(P)=\langle P,\mathcal{A}^{\top}_{\Omega}(P)\rangle$ and one easily verifies $\operatorname{grad}\widetilde{J}(P)=\operatorname{grad}J(P)$.
@@ 3.3 Action Functional
In [15] it has been shown that the assignment flow determined by (12) can be characterized as stationary point of a Lagrangian action functional, provided the affinity function $F$ satisfies a corresponding assumption. It turns out that this result applies to the patch assignment flow (20) which essentially has the same mathematical structure.
Proposition 2 (P-AF: action functional, stationary point)
@@ 5 Conclusion
Our further work will study in this context the design and structure of labeled patch dictionaries, as generators of nonlocal data labelings on graphs. In particular, discrete symmetries will be examined from the general viewpoint on locally equivariant networks that are generated by geometric flows, as developed by [7].
{credits}
5.0.1 Acknowledgements
This work is funded by the Deutsche Forschungsgemeinschaft (DFG) under Germany’s Excellence Strategy EXC-2181/1 - 390900948 (the Heidelberg STRUCTURES Excellence Cluster). This work is funded by the Deutsche Forschungsgemeinschaft (DFG), grant SCHN 457/17-2, within the priority programme SPP 2298: “Theoretical Foundations of Deep Learning”.

### v2

@@ Abstract
Keywords: assignment flows, Riemannian gradient flows, statistical manifolds, information geometry.
@@ 1 Introduction
Assignment flows [ÅPSS17] denote a class of dynamical systems evolving on statistical manifolds which serve as models for ‘neural ODEs’ on graphs obtained by geometric flow integration. A basic instance [SS21] are systems of the form $\dot{W}=R_{W}[\Omega W]$ (see Eq. 12 for the general form) which are Riemannian gradient flows with respect to a potential, parametrized by a spatial interaction matrix $\Omega(t)$, comprising a non-local graph Laplacian and a term steering the flow towards discrete decisions (label assignment). The time-variant parameters $\Omega(t)$ can be learned from data.
By vectorization (see Section 3.1 for details), this basic flow equation takes the form $\dot{w}=R^{\mathfrak{v}}_{w}[(\Omega\otimes I_{c})w]$ which suggests to generalize the parametrization to $\dot{w}=R^{\mathfrak{v}}_{w}[(\Omega\otimes\Omega_{c})w]$, in order to take additionally into account the interaction between $[c]=\{1,\dotsc,c\}$ labels, to be jointly assigned to data observed on a graph via the assignment flow $w(t)$. In fact, this formulation is general enough to cover a broad range of multi-population and multi-game dynamics [BCA+24].
@@
Related work. There is a vast literature on image denoising and restoration using continuous patches, either in a classical way by enforcing structure sparsity [CM12, Pey09, YSM12] or more recently by, e.g., Gaussian mixture priors [AAD+17, HBD18]. By contrast, our approach exploits labeled patches, where each label represents an equivalence class of continuous signals, for the structured prediction of labelings of metric data observed on a graph.
@@ 2 Preliminaries
We introduce few definitions from information geometry required in connection with assignment flows [AN00], [ÅPSS17], [AJLS17, Chapter 2]. Given a set $\mathcal{C}$ of category labels, we denote by $\mathcal{S}_{c}=\{w\in\mathbb{R}_{>0}:\langle\mathbb{1}_{c},w\rangle=1\}$ the relative interior of the probability simplex over $c=|\mathcal{C}|$ labels. This space forms a smooth manifold with trivial tangent bundle $T\mathcal{S}_{c}=\mathcal{S}_{c}\times T_{0}$ and tangent space at any $w\in\mathcal{S}_{c}$ given by $T_{0}:=T_{w}\mathcal{S}_{c}=\{v\in\mathbb{R}^{c}:\langle\mathbb{1}_{c},v\rangle=0\}$. $(\mathcal{S}_{c},g^{\mathcal{S}_{c}})$ becomes a Riemannian manifold when equipped with the Fisher-Rao metric
@@
Assignment flows [ÅPSS17] constitute an approach for determining a labeling function (2), based on the evolution of label assignment vectors as integral curves of vector fields on the assignment manifold, governed by a coupled system of replicator equations
@@
The function $F:\mathcal{S}_{c}^{n}\to\mathbb{R}^{n\times c}$ parametrizing the flow is often referred to as fitness (or payoff) function in evolutionary game theory and population dynamics, where analyses of individual (uncoupled) replicator equations have provided foundational insights into evolutionary stability and strategic dynamical games [HS98, San10]. For a more detailed exposition of assignment flows, we refer to [Sch20].
@@
Geometric integration of such flows can be performed numerically in a stable way using any of the methods worked out by [ZSPS20]. Under suitable conditions on the fitness function [ZZS22], these schemes guarantee convergence $\lim_{t\to\infty}W_{i}(t)=:e_{\ell_{\mathcal{V}}(i)},\;i\in\mathcal{V}$ to some labeling function (2), and they are amenable to learning parameters of the affinity function $F(W)$ from data [HSPS21, ZPS23, BGAPS24, CBP+24].
@@ 3.1 Flow Definition
Lemma 3.1 (maximizing patch consistency).
@@
Proof.  (16) results from (15) by taking into account the adjacency relation of $\mathcal{G}_{\mathcal{V}}$ given by $A^{h}_{\mathcal{V}}+A^{v}_{\mathcal{V}}=A_{\mathcal{V}}$, due to (1) and (7). The inner product with the first term on the right hand side of (15) reads
@@
Using the row-stacking vectorization map $\operatorname{vec}_{r}(ABC)=(A\otimes C^{\top})\operatorname{vec}_{r}(B)$ [VL00], we define the vectorized objective function
@@ 3.2 Independence of Graph Orientation
Proof. [sketch] Let $\widetilde{\mathcal{G}}_{\mathcal{V}}=(\mathcal{V},\widetilde{\mathcal{E}}_{\mathcal{V}})$ denote the transpose of $\mathcal{G}_{\mathcal{V}}$, that is the graph with the same vertex set $\mathcal{V}$ but with reversed edge orientations: $ij\in\mathcal{E}_{\mathcal{V}}\;\Leftrightarrow\;ji\in\widetilde{\mathcal{E}}_{\mathcal{V}},\;\text{for all }i,j\in\mathcal{V}$. Since the orientation affects the patch template adjacency matrix by (8), the resulting patch adjacency graph $\widetilde{\mathcal{G}}_{\mathcal{D}}$ becomes transposed as well. As a result, following Section 3.1, the corresponding objective then reads $\widetilde{J}(P)=\langle P,\mathcal{A}^{\top}_{\Omega}(P)\rangle$ and one easily verifies $\operatorname{grad}\widetilde{J}(P)=\operatorname{grad}J(P)$.
@@ 3.3 Action Functional
In [SAS24] it has been shown that the assignment flow determined by (12) can be characterized as stationary point of a Lagrangian action functional, provided the affinity function $F$ satisfies a corresponding assumption. It turns out that this result applies to the patch assignment flow (20) which essentially has the same mathematical structure.
Proposition 3.3 (P-AF: action functional, stationary point).
@@ 5 Conclusion
Our further work will study in this context the design and structure of labeled patch dictionaries, as generators of nonlocal data labelings on graphs. In particular, discrete symmetries will be examined from the general viewpoint on locally equivariant networks that are generated by geometric flows, as developed by [CSAS25].
Acknowledgments
This work is funded by the Deutsche Forschungsgemeinschaft (DFG) under Germany’s Excellence Strategy EXC-2181/2 - 390900948 (the Heidelberg STRUCTURES Excellence Cluster), and grant SCHN 457/17-2 within the priority programme SPP 2298: “Theoretical Foundations of Deep Learning.”